The 2 cordinates are: (4,-7) and (-7,4)
I need the:
Point Slope Form
Y-Interecept Form
A line parellel to this
A line perpendicular
Standard Form

Answer by Mathtut(3670)

first step is to find slope:
:
4-(-7)/-7-4=11/-11=-1.......so m=-1........now any line parallel will have the same slope and any line perpendicular will hava the negative inverse of -1
:
which is 1
:
so using one of our points y-4=-1(x+7)(point slope form--->y-4=-x-7
:
y=-x-3.....y intercept(slope intercept) form .
:
line parallel could be y=-x+5....in standard from would be x+y=5
:
line perpendicular could be y=x+1...in standard form would be x-y=-1